# Monster shocks, gamma-ray bursts and black hole quasi-normal modes from neutron-star collapse

### v1

@@ Abstract
We perform the first magnetohydrodynamic simulation tracking the magnetosphere of a collapsing magnetar. The collapse is expected for massive rotating magnetars formed in merger events, and it may occur many hours after the merger. The results suggest a novel gamma-ray burst (GRB) scenario, which creates a delayed high-energy counterpart of the merger gravitational waves. The collapse launches an outgoing magnetospheric shock, and a hot magnetized outflow forms behind the shock. The outflow is modulated by the ring-down of the nascent black hole, imprinting its kilohertz quasi-normal modes on the GRB tail.
@@
Introduction. Neutron star mergers are not only sources of gravitational waves, but are also accompanied by electromagnetic counterparts. Such counterparts provide insights into properties of dense matter in neutron stars and their magnetic fields. Canonical counterparts include a prompt gamma-ray burst (GRB [1]) and a kilonova ([2, 3]). Both were observed from the neutron star merger GW170817 (see, e.g., [4] for a summary). The GRB followed the merger with a 2-s delay and was likely emitted by a blast wave breaking out at the photosphere of the merger ejecta [5, 6, 7, 8]. The kilonova was emitted on a day timescale and powered by nuclear decay in the expanding ejecta.
@@
The explosion picture usually includes collimated relativistic jets from the black hole promptly formed after the merger. The jets are extremely bright sources of gamma-rays if observed face-on, and also emit a broad-band afterglow when they decelerate in an ambient medium [6]. However, some mergers may not promptly form a black hole and jets. Instead, they may form a long-lived neutron-star remnant [9, 10]. Its initial differential rotation is expected to generate magnetic fields up to $B=10^{16}$ G, which buoyantly emerge from the remnant and form its external magnetosphere [11, 12, 13, 14, 15, 16, 17, 18]. Then, the remnant cools and forms a young magnetar. It likely has a twisted magnetosphere, filled with $e^{\pm}$ pairs of a small mass density $\rho$. Its expected magnetization parameter $\sigma=B^{2}/\left(4\pi\rho c^{2}\right)$ is huge, exceeding $10^{10}$ [19].
@@
Eventually the remnant emits its angular momentum in a magnetized wind, loses its rotational support, and can collapse into a black hole [20, 21, 22]. The spindown occurs on the timescale $t_{\rm sd}\approx c^{3}I/\mu^{2}\Omega^{2}\approx 10^{4}\,\mu_{33}^{-2}(\nu/300\,{\rm Hz})^{-2}$ s, where $I\approx 10^{45}$ g cm2 is remnant’s moment of inertia, $\nu=\Omega/2\pi$ is its rotation rate and $\mu$ is the dipole moment of its magnetosphere (we normalized it to $10^{33}\,$ cm3). The delayed collapse after $t\sim t_{\rm sd}$ occurs suddenly, on a ms timescale.
@@
Previous simulations of neutron-star collapse followed the dynamics of the external magnetosphere either in vacuum or using force-free electrodynamics (FFE) [23, 24, 25]. Both frameworks do not allow plasma heating. Furthermore, both neglect plasma inertia and so are unable to track shock formation. By contrast, a full magnetohydrodynamic analysis predicts monster shocks, which can generate gamma-rays [19].
@@
This Letter reports the first magnetohydrodynamic simulations of the magnetosphere evolution in the dynamic spacetime of the collapsing star. It demonstrates shocks and ejection of a hot outflow that will emit a GRB. We also find that the outflow carries information about quasi-normal modes of the nascent black hole, which may be observed in the GRB time profile.
@@
Methods. We set up a simple initial state: a dipole magnetic field is attached to a rotating star with $\bm{\mu}\parallel\bm{\Omega}$. We artificially reduce the rotation rate from a realistic $\nu\gtrsim 300$ Hz to $56$  Hz, so that the light cylinder $R_{\rm LC}=c/\Omega$ is beyond our computational box. This simplification avoids the challenging preparation of an equilibrium rotating magnetosphere with the equatorial current sheet at $r>R_{\rm RL}$. Rotation is still essential, as it enables the ring-down effect when the star collapses into a black hole (e.g. [26]). The amplitude of ring-down oscillations scales with angular momentum. Therefore, we also simulate collapse with a high initial $\nu=900$ Hz; this additional simulation misses details of the outer magnetospheric dynamics but demonstrates the enhanced ring-down effect on the magnetized outflow from the nascent black hole.
@@
The star is initialized as an unstable general relativistic polytrope of mass $M=1.7M_{\odot}$ and radius $R_{\star}\approx 12$ km using the RNS code [27]; the details of its internal structure are unimportant, as we focus on the external magnetosphere. The magnetosphere has a low mass density $\rho$ and a large magnetization parameter $\sigma\gg 1$. At time $t=0$, we add a small pressure perturbation and the star begins to collapse.
@@
The simulation tracks the spacetime of the collapsing star, and evolves its magnetosphere according to general relativistic magnetohydrodynamics (GRMHD) equations [28]. We use the Frankfurt/IllinoisGRMHD (FIL) code [29, 30] which is built on top of the Einstein Toolkit infrastructure [31]. The spacetime dynamics is tracked using the Z4c formulation of the Einstein equations [32]. We use moving puncture coordinates [33] in the simulation and the presentation of results below. The GRMHD equations assume an ideal fluid with thermal pressure proportional to $e-\rho c^{2}$, where $e$ is the fluid energy density including rest mass. The equations are evolved using the ECHO scheme [34] with vector potential-based constraint transport [35, 36]. We employ a fixed three-dimensional Cartesian grid with $6$ levels of mesh refinement; the highest resolution has $78$ grid points per $R_{\star}$, and extend to $3.25\times R_{\star}$ in each direction.
@@
Unlike previous MHD simulations in dynamical spacetimes (e.g. [37, 38, 39, 40, 29]), our simulation follows the magnetosphere with a high magnetization parameter In particular, the initial “background” magnetosphere has $\sigma_{\rm bg}\sim 25$, and its perturbation during the collapse leads to ultra-relativistic motions with Lorentz factors $\gamma\sim 10$. Such simulations are challenging in terms of accuracy requirements and numerical stability of the algorithm, especially when MHD is coupled to a dynamically evolved spacetime. We have changed the code to improve its robustness (see also [41]), including different primitive recovery schemes [42, 43], drift floors [44], and bounds on $\sigma$ and $\beta=8\pi P/B^{2}$ where $P$ is the fluid pressure and $B$ is the magnetic field in fluid rest frame. Even with all these improvements, we have found that the fourth-order derivative corrector performs poorly in the shock region. Since the constraint-transport algorithm prevents us from switching it off selectively, we have disabled it everywhere. Thus, the simulation maintains second-order accuracy, different from all previous GRMHD simulations carried out with FIL [29, 45, 46, 12, 41].
@@
Collapse and wave launching. The collapsing star quickly forms a black hole with horizon radius $R_{h}\approx 2.5\,{\rm km}$ and the ergosphere around it. Effectively, the magnetospheric footprints on the star are quickly pulled in from $R_{\star}$ to $R_{h}$, and the star’s magnetic flux $\Psi$ is now threading the smaller sphere of radius $R_{h}$. As a result, a strong quasi-monopolar magnetic field $B\sim\Psi/r^{2}$ is created in the radial zone $R_{h}<r<R_{\star}$. This inner zone with the amplified magnetic pressure launches a compressive wave into the surrounding magnetosphere, which propagates with nearly speed of light, $v_{\rm wave}/c\approx 1-\sigma_{\rm bg}^{-1}$.
@@
Figure 1: Magnetohydrodynamic waves excited in the neutron star magnetosphere by the collapse. The rotation axis is along $z$, and the figure shows a cut of the magnetosphere along the plane of $y=0$. The black circle at the center is the nascent black hole. Green curves show the magnetic field lines, and color shows $E^{\phi}/B$ as measured by the local normal observer. The snapshot is taken at $t=0.67\,\rm ms$ from the onset of collapse.
@@
Monster shock formation. Magnetosonic waves were recently shown to accelerate plasma to a huge radial 4-velocity $u^{r}=\gamma v^{r}/c$ [19]. This effect occurs when the wave reaches radius $R_{\times}=(c\mu^{2}/8L)^{1/4}$ in the equatorial plane, where $\mu=r^{3}B_{\rm bg}$ and $L$ is the wave power. In the FFE limit ($\sigma_{\rm bg}\rightarrow\infty$) fluid expansion in the rarefaction phase of the compressive wave would diverge at $R_{\times}$ as $E^{2}-B^{2}$ touches zero and $u^{r}\rightarrow-\infty$. For a finite $\sigma_{\rm bg}$, fluid develops a finite $u^{r}\propto\sigma_{\rm bg}$. The ultrarelativistic fluid motion is directed toward the star and the wave immediately develops a monster shock.
Shock formation in magnetosonic waves has been demonstrated by kinetic plasma simulations [47] and by MHD calculations using characteristics [19]. For waves with frequency $\omega\gg c/R_{\times}$ a simple analytical MHD solution has been obtained. It demonstrates that at $r>R_{\times}$ the wave profile $E(t-r/c)$ develops a plateau of width $W_{p}\sim c/\omega$ where $E^{2}\approx B^{2}$. The plateau forms a linear accelerator, so the wave pushes the fluid 4-velocity to a huge value [19]
@@
The monster shock appears in our simulation where $E^{2}-B^{2}$ approaches zero, as predicted. We also observe the development of an $E$-plateau where $u^{r}$ develops a steep linear profile, reaching values consistent with Eq. (1) (see Fig. 2). The shock is the sudden jump of $u^{r}$ from a large negative value $u^{r}\sim-6$ back to a moderate $u^{r}$. These unique features of monster shocks are clear in the simulation despite numerical inaccuracies accumulated in the shock region. We also observe the expected strong heating localized at the shock (see the temperature panel in Fig. 2.)
The magnetization parameter $\sigma_{\rm bg}\sim 25$ used in the simulation is far below its real value in a magnetar, and the shock strength should be scaled to a larger $\sigma_{\rm bg}$ according to Eq. (1). The corresponding large $u^{r}\propto\sigma_{\rm bg}$ will make the shock highly radiative, i.e. the accelerated flow will radiate its energy before crossing the shock and joining the downstream flow [19]. Future simulations could attempt to track radiative transfer with self-consistent creation of $e^{\pm}$ pairs [48], which immediately make the flow optically thick. The dissipated energy is inevitably thermalized behind the shock, creating an opaque, radiation-dominated outflow. Our simulation assumes that the released energy remains trapped in the fluid. The observed relativistic outflow trails the shock, which expands with speed $c$.
@@
Black hole ring-down. The ring-down of the nascent spinning black hole lasts $\sim 100R_{h}/c$. It involves quasi-periodic oscillations of the horizon with frequencies characteristic of black hole quasi-normal modes, whose amplitude decays exponentially with time (e.g. [26]). For a stationary black hole, the quasi-monopolar magnetic flux $\Psi$ threading the horizon would be stuck for a significant time — its decay would be controlled by magnetic reconnection in the equatorial plane on a timescale $\gtrsim 100R_{h}/c$ [49]. By contrast, the oscillating black hole quickly and quasi-periodically sheds magnetic flux, losing significant $\delta\Psi$ each oscillation period. The discharged $\delta\Psi$ forms a quasi-periodic MHD outflow with the characteristic cusps of magnetic field lines in the equatorial plane. The cusps are inherited from the earlier split-monopole shape of field lines at the time of their decoupling from the oscillating black hole. This effect is seen in the simulation with the reduced rotation rate $\nu=56$ Hz (Fig. 1) and becomes stronger in the simulation with fast rotation $\nu=900$ Hz (Fig. 3).
Previous collapse simulations with a vacuum or FFE magnetosphere showed electromagnetic waves from ring-down [23, 24, 50, 25]. Vacuum ring-down is usually described as a coupling of black-hole quasi-normal modes to outgoing gravitational and electromagnetic waves [51, 52] using Newman-Penrose scalars

### v2

# Monster Shocks, Gamma-Ray Bursts, and Black Hole Quasi-normal Modes from Neutron-star Collapse

@@ Abstract
We perform the first magnetohydrodynamic simulation tracking the magnetosphere of a collapsing magnetar. The collapse is expected for massive rotating magnetars formed in merger events, and may occur many hours after the merger. Our simulation suggests a novel mechanism for a gamma-ray burst (GRB) which is uncollimated and forms a delayed high-energy counterpart of the merger gravitational waves. The simulation shows that the collapse launches an outgoing magnetospheric shock, and a hot magnetized outflow forms behind the shock. The outflow is baryon-free and uncollimated, and its power peaks on a millisecond timescale. Then, the outflow becomes modulated by the ring-down of the nascent black hole, imprinting its kilohertz quasi-normal modes on the GRB tail.
Black holes(162), General relativity(641), Gamma-ray bursts(629), High energy astrophysics(739), Neutron stars (1108), Plasma astrophysics(1261)
software: EinsteinToolkit (Loffler et al., 2012), Frankfurt/IllinoisGRMHD (Most et al., 2019; Etienne et al., 2015) RNS (Stergioulas & Friedman, 1995), kuibit (Bozzola, 2021), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), qnm (Stein, 2019), scipy (Virtanen et al., 2020)
1 Introduction
@@
Neutron star mergers are not only sources of gravitational waves, but are also accompanied by electromagnetic counterparts. Such counterparts provide insights into properties of dense matter in neutron stars and their magnetic fields. Canonical counterparts include a prompt gamma-ray burst (GRB (Meszaros, 2006)) and a kilonova (Li & Paczynski, 1998; Metzger, 2020). Both were observed from the neutron star merger GW170817 (see, e.g., Abbott et al. (2017) for a summary). The GRB followed the merger with a 2-s delay and was likely emitted by a blast wave breaking out at the photosphere of the merger ejecta (Murguia-Berthier et al., 2017; Gottlieb et al., 2018; Xie et al., 2018; Beloborodov et al., 2020). The kilonova was emitted on a day timescale and powered by nuclear decay in the expanding ejecta.
@@
The explosion picture usually includes collimated relativistic jets from the black hole promptly formed after the merger. The jets are extremely bright sources of gamma-rays if observed face-on, and also emit a broad-band afterglow when they decelerate in an ambient medium (Gottlieb et al., 2018). Similar jets are expected if the merger forms a short-lived neutron star with a debris accretion disk, see Rezzolla et al. (2011); Ruiz et al. (2016); Kawamura et al. (2016); Kiuchi et al. (2023) for magnetohydrodynamic (MHD) simulations of such systems. Such models predict the collapse of the neutron star in milliseconds to a second after the merger, triggered by accretion or the loss of differential rotation.
However, intermediate mass mergers may not promptly collapse to black holes and form jets (Bauswein et al., 2013; Köppel et al., 2019; Bauswein et al., 2020; Kashyap et al., 2022; Tootle et al., 2021; Kölsch et al., 2022; Schianchi et al., 2024). Instead, they may form a long-lived neutron-star remnant (Baiotti & Rezzolla, 2017; Radice et al., 2020). Its initial differential rotation is expected to generate magnetic fields up to $B=10^{16}$ G, which buoyantly emerge from the remnant and form its external magnetosphere (Kluzniak & Ruderman, 1998; Most & Quataert, 2023; Combi & Siegel, 2023; Giacomazzo & Perna, 2013; Giacomazzo et al., 2015; Kiuchi et al., 2015; Mösta et al., 2020; Aguilera-Miret et al., 2023). Then, the remnant cools and forms a young magnetar. It likely has a twisted magnetosphere, filled with $e^{\pm}$ pairs of a small mass density $\rho$. Its expected magnetization parameter $\sigma=B^{2}/\left(4\pi\rho c^{2}\right)$ is huge, exceeding $10^{10}$ (Beloborodov, 2023).
@@
Eventually, the remnant emits its angular momentum in a magnetized wind, loses its rotational support, and can collapse into a black hole (Lasky et al., 2014; Ravi & Lasky, 2014; Dall’Osso et al., 2015). The spindown occurs on the timescale $t_{\rm sd}\approx c^{3}I/\mu^{2}\Omega^{2}\approx 10^{4}\,\mu_{33}^{-2}(\nu/300\,{\rm Hz})^{-2}$ s, where $I\approx 10^{45}$ g cm2 is remnant’s moment of inertia, $\nu=\Omega/2\pi$ is its rotation rate and $\mu$ is the dipole moment of its magnetosphere (we normalized it to $10^{33}\,$ cm3). The delayed collapse after $t\sim t_{\rm sd}$ occurs suddenly, on a ms timescale. Proposed electromagnetic signals from the delayed collapse include low-frequency waves (Falcke & Rezzolla, 2014) and GRBs emitted by collimated jets (Ciolfi & Siegel, 2015).
@@
Previous simulations of neutron-star collapse followed the dynamics of the external magnetosphere either in vacuum or using force-free electrodynamics (FFE) (Baumgarte & Shapiro, 2003; Lehner et al., 2012; Dionysopoulou et al., 2013; Palenzuela, 2013; Most et al., 2018). Both frameworks do not allow plasma heating. Furthermore, both neglect plasma inertia and so are unable to track magnetospheric shock formation. By contrast, a full magnetohydrodynamic analysis predicts monster shocks, which can generate gamma-rays (Beloborodov, 2023).
@@
This Letter reports the first magnetohydrodynamic simulations of the magnetosphere evolution in the dynamic spacetime of the collapsing magnetar. It demonstrates shocks and ejection of a hot outflow that will emit a GRB. We also find that the outflow carries information about quasi-normal modes of the nascent black hole, which may be observed in the GRB time profile.
2 Methods
We set up a simple initial state: a dipole magnetic field is attached to a rotating star with $\bm{\mu}\parallel\bm{\Omega}$, using the vector potential given by Shibata et al. (2011) with a surface magnetic field of $B_{\ast}\sim 10^{16}\,\rm G$.
@@ 2 Methods
We artificially reduce the rotation rate from a realistic $\nu\gtrsim 300$ Hz to $56$  Hz, so that the light cylinder $R_{\rm LC}=c/\Omega$ is beyond our computational box. This simplification avoids the challenging preparation of an equilibrium rotating magnetosphere with the equatorial current sheet at $r>R_{\rm RL}$. Rotation is still essential, as it enables the ring-down effect when the star collapses into a black hole (e.g. Kokkotas & Schmidt (1999)). The amplitude of ring-down oscillations scales with angular momentum. Therefore, we also simulate collapse with a high initial $\nu=900$ Hz; this additional simulation misses details of the outer magnetospheric dynamics but demonstrates the enhanced ring-down effect on the magnetized outflow from the nascent black hole.
@@
We use the RNS code (Stergioulas & Friedman, 1995) to initialize the star as an unstable general relativistic polytrope (with polytropic coefficient K=100) of mass $M=1.7M_{\odot}$ and equatorial coordinate radius $R_{\star}\approx 12$ km; the details of its internal structure are unimportant, as we focus on the external magnetosphere. The magnetosphere has a low mass density $\rho$ and a large magnetization parameter $\sigma\gg 1$. At time $t=0$, we add a small pressure perturbation and the star begins to collapse.
The simulation tracks the spacetime of the collapsing star and evolves its magnetosphere according to general relativistic magnetohydrodynamics (GRMHD) equations (Duez et al., 2005). Since the plasma magnetosphere has a very small effective resistivity on the dynamical timescales of interests $t_{\rm dyn}\sim R/c$, it can be treated as an ideal conductor everywhere except the sites of magnetic reconnection, which develops later in the simulation (and is mediated by numerical resistivity).
@@
We use the Frankfurt/IllinoisGRMHD (FIL) code (Most et al., 2019; Etienne et al., 2015) which is built on top of the Einstein Toolkit infrastructure (Loffler et al., 2012). The spacetime dynamics is tracked using the Z4c formulation of the Einstein equations (Hilditch et al., 2013). We use moving puncture coordinates (Alcubierre et al., 2003) in the simulation and the presentation of results below. The GRMHD equations assume an ideal fluid with thermal pressure proportional to $e-\rho c^{2}$, where $e$ is the fluid energy density including rest mass. The equations are evolved using the ECHO scheme (Del Zanna et al., 2007) with vector potential-based constraint transport (Etienne et al., 2010, 2012).
No spatial symmetries are imposed during the simulation. We employ a fixed three-dimensional Cartesian grid with $6$ levels of mesh refinement; the highest resolution has $78$ grid points per $R_{\star}$. The grid extends to $3.25\times R_{\star}$ ($750\,\rm km$) in each direction. The boundaries of the computational domain are far from the shock and do not affect the results presented below.
@@
Unlike most MHD simulations in dynamical spacetimes (e.g. Liu et al. 2008; Kiuchi et al. 2014; Palenzuela et al. 2015; Ciolfi et al. 2017; Most et al. 2019), our simulation follows the magnetosphere with a high magnetization parameter (see also Paschalidis et al. (2015); Ruiz et al. (2016)). In particular, the initial “background” magnetosphere has $\sigma_{\rm bg}\sim 25$, and its perturbation during the collapse leads to ultra-relativistic motions with Lorentz factors $\gamma\sim 10$. Such simulations are challenging in terms of accuracy requirements and numerical stability of the algorithm, especially when MHD is coupled to a dynamically evolved spacetime.
We have changed the code to improve its robustness (see also Most (2023)), including different primitive recovery schemes (Kastaun et al., 2021; Kalinani et al., 2022), drift floors (Ressler et al., 2017), and bounds on $\sigma$ and $\beta=8\pi P/B^{2}$ where $P$ is the fluid pressure and $B$ is the magnetic field in fluid rest frame. Specifically, before black hole formation, we enforce bounds, $5010<\sigma<$ 50, and $\beta>0.02$, outside of the star. We distinguish between magnetospheric and stellar matter using a passive scalar (see also Ref. Parfrey & Tchekhovskoy (2017)). Even with all these improvements, we have found that the fourth-order derivative corrector performs poorly in the shock region. Since the constraint-transport algorithm prevents us from switching it off selectively, we have disabled it everywhere. Thus, the simulation maintains second-order accuracy, different from all previous GRMHD simulations carried out with FIL (Most et al., 2019, 2021; Chabanov et al., 2023; Most & Quataert, 2023; Most, 2023).
3 Results
@@
3.1 Collapse and wave launching
The collapsing star quickly forms a black hole with the apparent horizon radius $R_{h}\approx 2.5\,{\rm km}$ and the ergosphere around it. Effectively, the magnetospheric footprints on the star are quickly pulled in from $R_{\star}$ to $R_{h}$, and the star’s magnetic flux $\Psi$ is now threading the smaller sphere of radius $R_{h}$. As a result, a strong quasi-monopolar magnetic field $B\sim\Psi/r^{2}$ is created in the radial zone $R_{h}<r<R_{\star}$. This inner zone with the amplified magnetic pressure launches a compressive wave into the surrounding magnetosphere, which propagates with nearly speed of light, $v_{\rm wave}/c\approx 1-\sigma_{\rm bg}^{-1}$.
@@ 3.1 Collapse and wave launching
Figure 1: Magnetohydrodynamic waves excitepd in the neutron star magnetosphere by the collapse. The rotation axis is along $z$, and the figure shows a cut of the magnetosphere along the plane of $y=0$. The black circle at the center is the nascent black hole. Green curves show the magnetic field lines, and color shows $E^{\phi}/B$ as measured by the local normal observer. The snapshot is taken at $t=0.67\,\rm ms$ from the onset of collapse.
@@
3.2 Monster shock formation
Magnetosonic waves were recently shown to accelerate plasma to a huge radial 4-velocity $u^{r}=\gamma v^{r}/c$ (Beloborodov, 2023). This effect occurs when the wave reaches radius $R_{\times}=(c\mu^{2}/8L)^{1/4}$ in the equatorial plane, where $\mu=r^{3}B_{\rm bg}$ and $L$ is the wave power. In the FFE limit ($\sigma_{\rm bg}\rightarrow\infty$) fluid expansion in the rarefaction phase of the compressive wave would diverge at $R_{\times}$ as $E^{2}-B^{2}$ touches zero and $u^{r}\rightarrow-\infty$. For a finite $\sigma_{\rm bg}$, fluid develops a finite $u^{r}\propto\sigma_{\rm bg}$. The ultrarelativistic fluid motion is directed toward the star and the wave immediately develops a monster shock.
Shock formation in magnetosonic waves has been demonstrated by kinetic plasma simulations (Chen et al., 2022) and by MHD calculations using characteristics (Beloborodov, 2023). For waves with frequency $\omega\gg c/R_{\times}$ a simple analytical MHD solution has been obtained. It demonstrates that at $r>R_{\times}$ the wave profile $E(t-r/c)$ develops a plateau of width $W_{p}\sim c/\omega$ where $E^{2}\approx B^{2}$. The plateau forms a linear accelerator, so the wave pushes the fluid 4-velocity to a huge value (Beloborodov, 2023)
@@ 3.2 Monster shock formation
The monster shock appears in our simulation where $E^{2}-B^{2}$ approaches zero, as predicted. We also observe the development of an $E$-plateau where $u^{r}$ develops a steep linear profile, reaching values consistent with Eq. (1) (see Fig. 2). The shock is the sudden jump of $u^{r}$ from a large negative value $u^{r}\sim-6$ back to a moderate $u^{r}$. These unique features of monster shocks are clear in the simulation despite numerical inaccuracies accumulated in the shock region. We also observe the expected strong heating localized at the shock (see the temperature panel in Fig. 2).
The magnetization parameter $\sigma_{\rm bg}\sim 25$ used in the simulation is far below its real value in a magnetar, and the shock strength should be scaled to a larger $\sigma_{\rm bg}$ according to Eq. (1). The corresponding large $u^{r}\propto\sigma_{\rm bg}$ will make the shock highly radiative, i.e. the accelerated flow will radiate its energy before crossing the shock and joining the downstream flow (Beloborodov, 2023). Future simulations could attempt to track radiative transfer with self-consistent creation of $e^{\pm}$ pairs (Beloborodov, 2021), which immediately make the flow optically thick. The dissipated energy is inevitably thermalized behind the shock, creating an opaque, radiation-dominated outflow. Our simulation assumes that the released energy remains trapped in the fluid. The observed relativistic outflow trails the shock, which expands with speed $c$.
@@
3.3 Black hole ring-down
The ring-down of the nascent spinning black hole lasts $\sim 100R_{h}/c$. It involves quasi-periodic oscillations of the horizon with frequencies characteristic of black hole quasi-normal modes, whose amplitude decays exponentially with time (e.g. Kokkotas & Schmidt (1999)). For a stationary black hole, the quasi-monopolar magnetic flux $\Psi$ threading the horizon would be stuck for a significant time — its decay would be controlled by magnetic reconnection in the equatorial plane on a timescale $\gtrsim 100R_{h}/c$ (Bransgrove et al., 2021). By contrast, the oscillating black hole quickly and quasi-periodically sheds magnetic flux, losing significant $\delta\Psi$ each oscillation period. The discharged $\delta\Psi$ forms a quasi-periodic MHD outflow with the characteristic cusps of magnetic field lines in the equatorial plane. The cusps are inherited from the earlier split-monopole shape of field lines at the time of their decoupling from the oscillating black hole. This effect is seen in the simulation with the reduced rotation rate $\nu=56$ Hz (Fig. 1) and becomes stronger in the simulation with fast rotation $\nu=900$ Hz (Fig. 3).
Previous collapse simulations with a vacuum or FFE magnetosphere showed electromagnetic waves from ring-down (Baumgarte & Shapiro, 2003; Lehner et al., 2012; Palenzuela, 2013; Most et al., 2018). Vacuum ring-down is usually described as a coupling of black-hole quasi-normal modes to outgoing gravitational and electromagnetic waves (Teukolsky, 1972, 1973) using Newman-Penrose scalars